OpenStudy (anonymous):

Simplify (12x^-3 y^2)/(18xy^-1)Please help, i honestly have no idea how to do this

OpenStudy (shoba):

\[\frac{12 y ^{2} }{x ^{3} } divide \frac{ 18x }{ y } = \frac{12 y ^{2} }{x ^{3} } * \frac{ y}{ 18x }=\frac{ 12y ^{3} }{18x ^{4} } = \]

OpenStudy (shoba):

2y^3 / 3x^4

OpenStudy (anonymous):

@shoba thank u so much! i actually understand it now

OpenStudy (shoba):

no mention...have a nice day...